which is the equation of a parabola with vertex (0,0) and focus (0,2)?
Answer by lwsshak3(11628)

which is the equation of a parabola with vertex (0,0) and focus (0,2)?
**
Basic equation of a parabola that opens upward: (x-h)^2=4p(y-k), (h,k)=(x,y) coordinates of the vertex.
For given parabola:
vertex: (0,0)
axis of symmetry: x=0
p=2 (distance from vertex to focus on the axis of symmetry)
4p=8
equation of given parabola: x^2=8y